An electron has the following wave function:
¥ (x) = Ae-x (1-e -x)
Determine
i) the normalization constant A, and
ii) the expectation value of x
1
Expert's answer
2020-04-13T10:17:54-0400

The wave function is


"\u03c8(x)=Ae^{-x}(1-e^{ -x})."

Determine the normalization constant A:


"1=\int^{+\infty}_{-\infty}\psi^*(x)\psi(x)dx,\\
\space\\
1=\int^{+\infty}_{0}A^2e^{-2x}(1-e^{-x})2\space dx,\\
\space\\
1=A^2\bigg(\frac{1}{12}\bigg),\\
A^2=12.\space\\"

Determine the expectation value of x:


"<\hat{A}>=\int^{+\infty}_{-\infty}\psi^*(x)\hat{A}\psi(x)dx,\\
\space\\
<x>=12\int^{+\infty}_{0}xe^{-2x}(1-e^{-x})2\space dx,\\
\space\\<x>=\frac{13}{12}."